CHM 116: General Chemistry 2, Fall 2018 Ch 17 Assessment em 5 Constants 1 Perlodic Table Part A A 1.0-L bufer solution is 0.105 d in HNO2 and 0.185 Min NaNO Determine the concentrations of HNO, and NaNO ater addition of 1.5 g HCI Express your answers using three significant figures separated by a comma. IRNO,. NavO,-0.146,1.609 X Incorrect; Try Again; 3 attempts remaining Part B ooncentrations of HNO2 and NaNO2 atar addition of 1 5 g NaOH Express your answers using three significent figures separated by a comma 名
Answer #1

A) We determine the reaction that will occur with HCl:

NaNO2 + HCl = HNO2 + HCl

We determine the moles of HCl added:

Mol HCl = 1.5 gr * (1 mol / 36.46 gr) = 0.041 mol

Then 0.041 mol is subtracted from NaNO2 and 0.041 mol is added to HNO2:

[NaNO2] = 0.165 - 0.041 = 0.124 M

[HNO2] = 0.105 + 0.041 = 0.146 M

B) Similar methodology: Reaction:

HNO2 + NaOH = NaNO2 + H2O

moles added:

Mol NaOH = 1.5 gr (1 mol / 40 gr) = 0.038 mol

Then to NaNO2 add 0.038 mol and to HNO2 subtract 0.038 mol:

[NaNO2] = 0.165 + 0.038 = 0.203 M

[HNO2] = 0.105 - 0.038 = 0.067 M.